OpenStudy (anonymous):

Find the vertex and axis of symmetry of the graph of the function. f(x) = x^2 + 13x + 6

OpenStudy (ybarrap):

x^2 + 13x + 6 (x + 13/2)^2 - (13/2)^2 + 6 = (x + 13/2)^2 - (13/2)^2 + 6 Vertex is minimum when x = - 13/2 This is also the axis of symmetry